28 tea pluckers can pluck 1260 kg of leaves in 8 hours. how many more people will be required, if 1200 kg of tea need to be plucked in 6 hours?
Answer by addingup(3677)

a rate-of-work problem.
28*8 = 224 man-hours = 224*60 = 13440 minutes
13440/1200 = 11.2 man-minutes per kg
------------------------------------------
360/11.2 = 32.14